(13.36) lere is the distribution of the adjusted gross income (in thousands of dollars) reported on individual federal income tax returns in 2011: 15- 29 30 71 0.315 Income 15 75 199 200 Probablity 0.264 0.216 0. 179 026 (a) What is the probability that a randomly chosen return shows an adjusted gross incorme of $30,000 or more? P(income 30,0) +0.00)- (b) Given that a retum shows an Income of at least $30,000, what is the conditional prohabllity that the Income is at least $75,000 P (income $75,000 income S 30,000) (40.001)
Answer #1

a) P(income >=30000)=1-0.264-0.216=0.52

b) income at least 30000= 0.52
income at least 75000= 1-0.264-0.216-0.315=0.205
required probability =0.205/0.52

=0.3942

.............................

if you have doubt ask in comment give thumbs up if you like work.